Six artificial satellites complete one circular or
Answer #1

ω is the same for all satellites.

T = ω2 m L, scaling as the product of m and L

a) ml = 800*5000 = 4*10^6

b) ml = 100*2500 = 0.25*10^6

c) ml = 100*10000 = 1*10^6

d) ml = 200*5000 = 1*10^6

e) ml = 300*7500 = 2.25*10^6

f) ml = 400*2500 = 1*10^6

a,e,c=d=f,b